OpenStudy (anonymous):

A store is having a sale on trail mix and jelly beans. For 7 pounds of trail mix and pounds of jelly beans, the total cost is $53 . For 5 pounds of trail mix and 3 pounds of jelly beans, the total cost is $25 . Find the cost for each pound of trail mix and each pound of jelly beans.

OpenStudy (wolfe8):

Can you check your second sentence? I think you missed a number

OpenStudy (anonymous):

9

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

9 pounds of jelly beans

OpenStudy (wolfe8):

Alright so can you put what info you have into mathematical equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if you could do that i might be able to solve it myself

OpenStudy (wolfe8):

Okay let's work together. We have "7 pounds of trail mix and 9 pounds of jelly beans, the total cost is $53". Let x be trail mix and y be jelly beans. It then translates to 7x+9y=53 Can you try for the next line of info?

OpenStudy (anonymous):

5x+3y=25

OpenStudy (wolfe8):

Correct. And can you solve it now?

OpenStudy (anonymous):

each one ?

OpenStudy (wolfe8):

For x and y, which are prices for trail mix and jelly beans per pound respectively. Do you know how to solve simultaneous equations?

OpenStudy (anonymous):

subtitution method?

OpenStudy (wolfe8):

Yes and a few other methods, but you can do that definitely. Go ahead and show me what you get.

OpenStudy (anonymous):

i solve for y for the first one?

OpenStudy (wolfe8):

You can. It actually doesn't matter what you solve for as the first one. Just look out for the easiest one if you can(i.e the variable with coefficient 1)

OpenStudy (anonymous):

y=37/15 x=22/5

OpenStudy (wolfe8):

That doesn't look quite right to me. Can you show me the working?

OpenStudy (anonymous):

right x=11/4 y=15/4

OpenStudy (wolfe8):

That looks about right. The way to check is to put in those values in any equation and see if it's the same as the right hand side.

OpenStudy (anonymous):

then

OpenStudy (wolfe8):

Well price of trail mix per pound is $11/4 per pound and jelly beans $15/4 per pound. But people don't usually write money in that form so get the decimal form.

OpenStudy (anonymous):

how?

OpenStudy (wolfe8):

Fractions also mean the numerator divided by the denominator. Does that help?

OpenStudy (anonymous):

yea 2.75

OpenStudy (wolfe8):

Yes that is for mix trail. Do the same for jelly beans and you have your answers.

OpenStudy (anonymous):

thanks

OpenStudy (wolfe8):

You're welcome. Happy studying!